Will a precipitate form if 120 mL of 0.045 M silver nitrate is mixed with 180 mL of 0.052 M sodium
chromate? (silver chromate has Ksp = 1.9 x10^–12) Calculate [Ag+] and [CrO4^2–] at equilibrium.
1
Expert's answer
2020-10-30T06:44:56-0400

2Ag+ (aq) + CrO42- (aq) ↔ Ag2CrO4 (s)

mol Ag+ (aq) = (0.120 L)(0.045 mol/L) = 0.0054 mol Ag+

[Ag+] = 0.0054 mol/(0.120 L + 0.180 L) = 0.0054 mol / 0.300 L = 0.018 M

mol CrO42- (aq) = (0.180 L)(0.052 mol/L) = 0.00936 mol CrO42-

[CrO42-] = 0.00936 mol / 0.300 L = 0.0312 M

Q = [Ag+]2[CrO42-] = (0.018)2(0.0312) = 1.01 x 10-5

Q > Ksp

Ag2CrO4 will precipitate from solution.

At the point where Q = Ksp the solution is at equilibrium.

Assumption: [Ag+] = [CrO42-]

1.9 x 10-12 = [Ag+]2[Ag+] = [Ag+]3

[Ag+] = [CrO42-] = 1.24 x 10-4 M (at equilibrium)